. Let f and g be differentiable functions satisfying g'(a) = 2, g(a) = b, and fog = 1 (identity function). Then f'(b) is equal to
A. 0 B. 2/3 C. 1/2 D. none of the above

Grade:Upto college level

2 Answers

Saurabh Koranglekar
3 years ago
Dear student

The answer is option c

Regards
Vikas TU
3 years ago
fog x = f(g(x)) 
f'(g(x)) * g'(x) = 1 
f'(g(x)) = 1/g'(x) 
g(x) = b , x = a 
f'(b) = 1/g'(a) = 1/2 
….....................................